# A Dynamic, Self-balancing k-d Tree

### v11

@@ 1 Introduction
Bentley introduced the k-d tree as a binary search tree that stores k-dimensional data [Bentley]. Like a binary search tree, a k-d tree partitions a set of data at each recursive level of the tree. Unlike a binary search tree that uses only one key for all levels of the tree, a k-d tree uses $k$ keys and cycles through the keys for successive levels of the tree. For example, to build a k-d tree from three-dimensional tuples that represent $\left(x,y,z\right)$ coordinates, the keys would be cycled as $x,y,z,x,y,z$... for successive levels of the k-d tree. A more elaborate scheme for cycling the keys chooses the coordinate that has the widest dispersion or largest variance to be the key for a particular level of recursion [Friedman], but this article assumes that the keys are cycled as $x,y,z,x,y,z$... in order to simplify the exposition.
@@ 1.1 Two Definitions of Balance
The balance at each node of a dynamic k-d tree is computed from the heights of the node’s left and right subtrees, where height is defined as the maximum path length from the node to the bottom of the tree. The height of a null child pointer is 0. The height of a leaf node is 1. The height of any particular node is computed as 1 plus the greater of the heights of its two children, and is assigned to that node’s height field.
@@ 1.2 Insertion into a k-d Tree and Subsequent Rebalancing
Then, as the recursion unwinds, the height is computed at each node along the return path to the root of the tree, and the balance is checked at each node to determine whether rebalancing is required at that node. In Figure 1, the difference in child subtree heights is 0, 1, and 2 at nodes $\left(9,2,1\right)$; $\left(8,3,2\right)$; and $\left(8,1,5\right)$ respectively. Hence, the subtree rooted at node $\left(8,1,5\right)$ must be rebalanced, as required by the AVL or red-black criterion because either criterion enforces a subtree height difference of at most one node. This subtree is rebalanced by rebuilding it [Overmars] to create the subtree shown in Figure 2 that contains nodes $\left(9,2,1\right)$; $\left(8,3,2\right)$; and $\left(8,1,5\right)$ depicted with dashed borders, and whose respective $z$:$x$:$y$ super keys have relative order 11:99:22 << 22:88:33 << 55:88:11. Rebuilding a subtree that contains 3 nodes or fewer, such as this subtree, occurs frequently and requires comparison of at most 3 super keys [Stepanov]. Rebuilding a subtree that contains more than 3 nodes requires a more elaborate tree-building algorithm that has, for example, computational complexity $O\left[kn\log\left(n\right)\right]$ or $O\left[n\log\left(n\right)\right]$ [Brown2015].
@@ 1.3 Deletion from a k-d Tree and Subsequent Rebalancing
The first case is deletion of a leaf node that has no children. The node is removed from the tree. Then the recursion unwinds, with computation of the height and potential rebalancing at each node along the return path to the root of the tree.
@@
However, finding the predecessor or successor node is more complicated for a k-d tree than for a 1-d tree [Drozdek] [Rocca]. The predecessor node to a particular node in a 1-d tree is found by beginning with the less-than child to that node, and recursively following the greater-than child pointer until it is null. The node with the null greater-than pointer is the predecessor node. The successor node to a particular node is found in a similar manner by beginning with the greater-than child to that node, and recursively following the less-than child pointer until it is null.
@@
Figure 3: For finding the predecessor node in a k-d tree, solid arrows designate child pointers that must be followed recursively, and solid borders designate potential predecessor nodes.
@@
Because node $b$ has a $y$:$z$:$x$ super key, its $x$:$y$:$z$ sorted order relative to node $a$ is unknown, so it must inspected as a potential predecessor node to node $a$. Also, because of its $y$:$z$:$x$ super key, node $b$ cannot eliminate its children (nodes $c$ and $d$) from recursive search for the predecessor node to node $a$ that has an $x$:$y$:$z$ super key.
Nodes $c$ and $d$ have $z$:$x$:$y$ super keys, so their $x$:$y$:$z$ sorted orders relative to node $a$ are unknown, and hence they must be inspected as potential predecessor nodes to node $a$. Also, because of their $y$:$z$:$x$ super keys, nodes $c$ and $d$ cannot eliminate their children (nodes $e$, $f$, $g$ and $h$) from recursive search for the predecessor node.
Nodes $e$, $f$, $g$ and $h$ have $x$:$y$:$z$ super keys, so their $x$:$y$:$z$ sorted orders relative to node $a$ are known. In that case, the rules for selection of a potential predecessor node are similar to the rules for selection of a predecessor node in a 1-$d$ tree. Specifically, if any of nodes $e$, $f$, $g$ and $h$ have a null greater-than child pointer, those nodes ($f$ and $h$) are potential predecessor nodes. Conversely, if any of nodes $e$, $f$, $g$ and $h$ do not have a null greater-than child pointer, those nodes ($e$ and $g$) cannot be potential predecessor nodes, so their greater-than child pointers must be followed recursively to identify potential predecessor nodes. But it is not necessary to follow their less-than child pointers recursively, because no node in a less-than branch can have the largest $x$:$y$:$z$ super key.
In view of this exclusion of all less-than branches, finding a 1-child node’s immediate predecessor node requires inspecting $n^{1-1/k}$ nodes of the subtree rooted at node $b$, assuming that $n$-node subtree to be perfectly balanced [Drozdek], [Rocca] [Samet]. And finding a 1-child node’s immediate successor node requires similarly searching the subtree rooted at that 1-child node’s greater-than child, to find the node that has the smallest $x$:$y$:$z$ super key.
When the immediate predecessor or immediate successor node to the 1-child node has been found in the manner described above, that predecessor or successor node’s tuple field is copied to the 1-child node’s tuple field, and then the predecessor or successor node is deleted from the tree. But that predecessor or successor node is not necessarily a leaf node, and if not, it must be deleted recursively as one of the three cases itemized above [Rocca], [Samet]. Recursive deletion converges to a leaf node that is removed from the tree. Then the recursion unwinds, with computation of the height and potential rebalancing at each node along the return path.
@@
Recursive deletion of a non-leaf replacement node converges to a leaf node that is removed from the tree. Then the recursion unwinds, with computation of the height and potential rebalancing at each node along the return path to the root of the tree.
@@ 1.4 Benchmark Methodology
Benchmarks for the k-d tree were executed for trees that contained $n$ nodes, for values of $n$ in the range $\left[1,003,201;4,523,071\right]$ that map to equally spaced values of $n\log_{2}\left(n\right)$ in the range $\left[20,000,000;100,000,000\right]$. Each node of the tree stored a $k$-dimensional tuple of 64-bit integers. The integers were equally spaced across the maximum 64-bit integer range $r$, so the spacing was $r/n$. The integers were randomly shuffled via the std::mt19937_64 Mersenne Twister pseudo-random number generator [Matsumoto] and copied to the first of the $k$ dimensions, then randomly shuffled again and copied to the second of the $k$ dimensions, et cetera.
So that all benchmarks randomly shuffled the integers in an identical sequence, each benchmark initialized std::mt19937_64 to std::mt19937_64::default_seed.
The benchmark for the k-d tree was implemented in C++, compiled via Gnu g++ 13.2.0 with the -std=c++20, -O3, -pthread -D NLOGN -D ENABLE_PREFERRED_TEST options that specify red-black balance, and executed under Ubuntu 24.04.1 LTS via a single thread mapped to a single performance core via the Ubuntu taskset command.
Each benchmark measured the execution times for insertion of all tuples into the tree, verification of correct ordering of the tree, search for all tuples in the tree, and deletion of all tuples from the tree via the std::chrono::steady_clock::now() function. To provide a comparison, the execution time was measured for construction of a static k-d tree via an $O\left[n\log\left(n\right)\right]$ algorithm [Brown2015]. Each benchmark was repeated 100 times and the mean values and standard deviations of the execution times were calculated. The standard deviations were less than 1% of the mean values.
@@ 1.5 Benchmark Results
Figure 4 plots execution times (in seconds) for insertion into, deletion from, and search of a dynamic k-d tree, plotted versus $n\log_{2}\left(n\right)$ where $n$ represents the number of tuples in the tree. Execution times for random and sorted (worst-case) tuples are included. Also included are execution times for construction of a static k-d tree.
@@ 1.6 Choosing the Replacement Node for a Deleted Two-child Node
Section 1.3 explains that a deleted 2-child node may be replaced by either its immediate predecessor node or its immediate successor node, and that the replacement node could be obtained from the higher of the deleted node’s child subtrees. Table 2 reports execution times (in seconds) for deletion of random tuples from a k-d tree, wherein the replacement node for a deleted 2-child node is obtained either from the higher child subtree or always from the greater-than subtree (i.e., the immediate successor node). This table reveals that neither node replacement strategy is consistently faster than the other, and the difference between deletion execution times is 1% or less.
@@ 1.7 AVL Versus Red-black Balance
As explained in Section 1.1 of this article, the balance of a k-d tree may be established either via the red-black criterion, which requires that the heights of the two child subtrees of a node differ by at most a factor of two; or via the AVL criterion, which requires that the heights of the two child subtrees of a node differ by 1. In fact, it is possible to define the AVL criterion less stringently; for example, to allow the subtrees to differ by an integer in the range $\left[1,4\right]$ instead of by 1 [Foster1973].
@@ 2 Conclusions
A dynamic k-d tree confers flexibility to building a k-d tree. Instead of building a static, balanced k-d tree en masse from all of the k-dimensional data, it is possible to build the tree piecemeal, one datum at a time. Alternately, it is possible to build a static, balanced tree from the available data, and thereafter modify the tree dynamically via insertion or deletion, one datum at a time. For either approach to building a tree, the dynamic k-d tree algorithms ensure that the tree remains balanced.
The algorithm that inserts a k-dimensional tuple into a dynamic k-d tree is similar to the algorithm that inserts a 1-dimensional key into a standard binary search tree. The algorithm that deletes a k-dimensional tuple from a dynamic k-d tree differs from the algorithm that deletes a 1-dimensional key from a standard binary search tree.
@@
Included with this manuscript are Java and C++ implementations of $O\left[n\log\left(n\right)\right]$ algorithms that insert into and delete from a dynamic k-d tree or a dynamic k-d tree-based key-to-multiple-values map, as well as $O\left[n\log\left(n\right)\right]$ and $O\left[kn\log\left(n\right)\right]$ algorithms that build a static k-d tree or a static k-d tree-based key-to-multiple-values map and that rebalance a subtree of the dynamic tree or map.
The dynamic k-d tree is implemented as the KdTreeDynamic class that derives from or extends the KdTree class that implements the $O\left[n\log\left(n\right)\right]$ and $O\left[kn\log\left(n\right)\right]$ algorithms that build a static k-d tree. Hence, all other algorithms implemented by the KdTree class, such as region search, nearest-neighbor search and reverse-nearest-neighbor search [Korn], are accessible by the KdTreeDynamic class. In addition, it is possible to build a static k-d tree via an $O\left[n\log\left(n\right)\right]$ or $O\left[kn\log\left(n\right)\right]$ algorithm, and then provide that static tree to the KdTreeDynamic constructor so that the static tree may thereafter be updated dynamically via insertion or deletion of individual tuples.
The Java and C++ implementations may be downloaded at the following URL.

### v12

@@ 1 Introduction
Bentley introduced the k-d tree as a binary search tree that stores k-dimensional data [Bentley]. Like a binary search tree, a k-d tree partitions a set of data at each recursive level of the tree. Unlike a binary search tree that uses only one key for all levels of the tree, a k-d tree uses $k$ keys and cycles through the keys for successive levels of the tree. For example, to build a k-d tree from three-dimensional tuples that represent $\left(x,y,z\right)$ coordinates, the keys would be cycled as $x,y,z,x,y,z$... for successive levels of the k-d tree. A more elaborate scheme for cycling the keys chooses the coordinate that has the widest dispersion or largest variance to be the key for a particular level of recursion [Friedman] but this article assumes that the keys are cycled as $x,y,z,x,y,z$... in order to simplify the exposition.
@@ 1.1 Two Definitions of Balance
The balance at each node of a dynamic k-d tree is computed from the heights of the node’s left and right subtrees, where height is defined as the maximum path length from the node to the bottom of the tree. The height of a null child pointer is 0. The height of a leaf node is 1, and the height of a non-leaf node is 1 plus the greater of the heights of its two children. A node’s height is assigned to that node’s height field.
@@ 1.2 Insertion into a k-d Tree and Subsequent Rebalancing
Then, as the recursion unwinds, the height is computed at each node along the return path to the root of the tree, and the balance is checked at each node to determine whether rebalancing is required at that node. In Figure 1, the difference in child subtree heights is 0, 1, and 2 at nodes $\left(9,2,1\right)$; $\left(8,3,2\right)$; and $\left(8,1,5\right)$ respectively. Hence, the subtree rooted at node $\left(8,1,5\right)$ must be rebalanced, as required by the AVL or red-black criterion because either criterion enforces a subtree height difference of at most one node. This subtree is rebalanced by rebuilding it [Overmars] to create the subtree shown in Figure 2 that contains nodes $\left(9,2,1\right)$; $\left(8,3,2\right)$; and $\left(8,1,5\right)$ depicted with dashed borders, and whose respective $z$:$x$:$y$ super keys have relative order 11:99:22 << 22:88:33 << 55:88:11. Rebuilding a subtree that contains 3 nodes or fewer, such as this subtree, occurs frequently and requires comparison of at most 3 super keys [Stepanov]. Rebuilding a subtree that contains more than 3 nodes requires a comprehensive tree-building algorithm, such as a $O\left[kn\log\left(n\right)\right]$ or $O\left[n\log\left(n\right)\right]$ parallel algorithm [Brown2015].
@@ 1.3 Deletion from a k-d Tree and Subsequent Rebalancing
The first case is deletion of a leaf node that has no children. The leaf node is excised from the tree. Then the recursion unwinds, with computation of the height and potential rebalancing at each node along the return path to the root of the tree.
@@
However, finding the predecessor or successor node is more complicated for a k-d tree than for a 1-d tree [Drozdek] [Rocca]. The predecessor node to a particular node in a 1-d tree is found by starting with the less-than child to that node, and recursively following the greater-than child pointer until it is null, because the null-pointer node has the largest key in the less-than subtree. The successor node to a particular node is found in a similar manner by starting with the greater-than child to that node, and recursively following the less-than child pointer until it is null.
@@
Figure 3: Finding the predecessor node in a k-d tree: solid arrows designate child pointers that must be followed recursively, and solid borders designate potential predecessor nodes.
@@
Because node $b$ has a $y$:$z$:$x$ super key, its $x$:$y$:$z$ sorted order relative to node $a$ is unknown, so it must inspected as a potential predecessor node to node $a$. Also, because of its $y$:$z$:$x$ super key, node $b$ cannot eliminate its children (nodes $c$ and $d$) from recursive search for node $a$’s predecessor node.
Because nodes $c$ and $d$ have $z$:$x$:$y$ super keys, their $x$:$y$:$z$ sorted orders relative to node $a$ are unknown, so they must be inspected as potential predecessor nodes to node $a$. Also, because of their $z$:$x$:$y$ super keys, nodes $c$ and $d$ cannot eliminate their children (nodes $e$, $f$, $g$ and $h$) from recursive search for node $a$’s predecessor node.
Nodes $e$, $f$, $g$ and $h$ have $x$:$y$:$z$ super keys, so their $x$:$y$:$z$ sorted orders relative to node $a$ are known. In that case, selection of a potential predecessor node is similar to selection of the predecessor node in a 1-$d$ tree. Specifically, if any of nodes $e$, $f$, $g$ and $h$ have a null greater-than child pointer, those nodes ($f$ and $h$) are potential predecessor nodes. But if any of nodes $e$, $f$, $g$ and $h$ do not have a null greater-than child pointer, those nodes ($e$ and $g$) cannot be potential predecessor nodes; hence, their greater-than child pointers must be followed recursively to inspect potential predecessor nodes $i$ and $l$ respectively. In neither case is it necessary to follow recursively the less-than child pointers of nodes $f$ and $g$ to inspect nodes $j$ and $k$ respectively, because no node in these less-than branches can have the largest $x$:$y$:$z$ super key.
In view of the exclusion of less-than branches, finding 1-child node $a$’s immediate predecessor node requires inspecting $n^{1-1/k}$ nodes of the $n$-node subtree rooted at node $b$, assuming that subtree to be perfectly balanced with no empty branches [Samet]. And if 1-child node $a$ has only a greater-than child instead of a less-than child, finding node $a$’s immediate successor node requires searching the subtree rooted at that greater-than child to find the node that has the smallest $x$:$y$:$z$ super key.
When the immediate predecessor or immediate successor node to the 1-child node has been found in the manner described above, that predecessor or successor node’s tuple field is copied to the 1-child node’s tuple field, and then the predecessor or successor node is deleted from the tree. But that predecessor or successor node is not necessarily a leaf node, and if not, it must be deleted recursively as one of the three cases itemized above [Rocca], [Samet]. Recursive deletion converges to a leaf node that is excised from the tree. Then the recursion unwinds, with computation of the height and potential rebalancing at each node along the return path.
@@
Recursive deletion of a non-leaf replacement node always converges to a leaf node that is excised from the tree. Then the recursion unwinds while computing the height and potentially rebalancing at each node along the return path to the root of the tree.
@@ 1.4 Benchmark Methodology
A k-d tree benchmark measured algorithm performance for trees that contained $n$ nodes, for values of $n$ in the range $\left[1,003,201;4,523,071\right]$ that map to equally spaced values of $n\log_{2}\left(n\right)$ in the range $\left[20,000,000;100,000,000\right]$. Each node of the tree stored a $k$-dimensional tuple of 64-bit integers. The integers were equally spaced across the maximum 64-bit integer range $r$, so the spacing was $r/n$. The integers were randomly shuffled via the std::mt19937_64 Mersenne Twister pseudo-random number generator [Matsumoto] and copied to the first of the $k$ dimensions, then randomly shuffled again and copied to the second of the $k$ dimensions, et cetera. So that all benchmarks randomly shuffled the integers in an identical sequence, each benchmark initialized std::mt19937_64 to std::mt19937_64::default_seed.
The k-d tree benchmark was implemented in C++, compiled via Gnu g++ 13.2.0 with -std=c++20, -O3, -pthread -D NLOGN options to specify red-black balance and a $O\left[n\log\left(n\right)\right]$ rebuilding algorithm, and executed under Ubuntu 24.04.1 LTS via a single thread mapped to a single performance core via the Ubuntu taskset command.
Each benchmark measured the execution times for insertion of all tuples into the tree, verification of correct ordering of the tree, search for all tuples in the tree, and deletion of all tuples from the tree via the std::chrono::steady_clock::now() function. To provide a comparison, the execution time was measured for construction of a static k-d tree via a $O\left[n\log\left(n\right)\right]$ algorithm. Each benchmark was repeated 100 times and the mean values and standard deviations of the execution times were calculated. The standard deviations were less than 1% of the mean values.
@@ 1.5 Benchmark Results
Figure 4 plots execution times (in seconds) for single-threaded insertion into, deletion from, and search of a dynamic k-d tree, plotted versus $n\log_{2}\left(n\right)$ where $n$ represents the number of tuples in the tree. Execution times are included for random and sorted (worst-case) tuples, as well as for single-threaded construction of a static k-d tree.
@@ 1.6 Choosing the Replacement Node for a Deleted Two-child Node
Section 1.3 proposes that a deleted 2-child node may be replaced by either its immediate predecessor node or its immediate successor node, and that the replacement node may be obtained from the higher of the deleted node’s child subtrees. Table 2 reports execution times (in seconds) for deletion of random tuples from a k-d tree, wherein the replacement node for a deleted 2-child node is obtained either from the higher child subtree or always from the greater-than subtree (i.e., the immediate successor node). This table reveals that neither node replacement strategy is consistently faster than the other, and the difference between deletion execution times is 1% or less.
@@ 1.7 AVL Versus Red-black Balance
Section 1.1 of this article explains that the balance of a k-d tree may be established either via the red-black criterion, which requires that the heights of the two child subtrees of a node differ by at most a factor of two; or via the AVL criterion, which requires that the heights of the two child subtrees of a node differ by 1. In fact, it is possible to define the AVL criterion less stringently; for example, to allow the subtrees to differ by an integer in the range $\left[1,4\right]$ instead of by 1 [Foster1973].
@@ 2 Conclusions
A dynamic k-d tree confers flexibility to building a k-d tree. Instead of building a static, balanced k-d tree en masse from all of the k-dimensional data, it is possible to build the tree piecemeal, one datum at a time. Alternately, it is possible to build a static, balanced tree from the available data, and thereafter modify the tree dynamically via insertion or deletion, one datum at a time. For either approach to building a tree, dynamic k-d tree algorithms ensure that the tree remains balanced.
The algorithm that inserts a k-dimensional tuple into a dynamic k-d tree is similar to the algorithm that inserts a 1-dimensional key into a standard binary search tree. The algorithm that deletes a k-dimensional tuple from a dynamic k-d tree is more complicated than the algorithm that deletes a 1-dimensional key from a standard binary search tree.
@@
Included with this manuscript are Java and C++ implementations of $O\left[n\log\left(n\right)\right]$ algorithms that insert into and delete from a dynamic k-d tree or a dynamic k-d tree-based key-to-multiple-values map. Also included are $O\left[n\log\left(n\right)\right]$ and $O\left[kn\log\left(n\right)\right]$ algorithms that build a static k-d tree or a static k-d tree-based key-to-multiple-values map and that rebalance a subtree of the dynamic tree or map.
The dynamic k-d tree is implemented as the KdTreeDynamic class that derives from (in C++), or extends (in Java), the KdTree class that implements the $O\left[n\log\left(n\right)\right]$ and $O\left[kn\log\left(n\right)\right]$ algorithms that build a static k-d tree. Hence, all other algorithms implemented by the KdTree class, such as region search, nearest-neighbor search and reverse-nearest-neighbor search [Korn], are accessible by the KdTreeDynamic class. In addition, it is possible to build a static k-d tree via an $O\left[n\log\left(n\right)\right]$ or $O\left[kn\log\left(n\right)\right]$ algorithm, and then provide that static tree to the KdTreeDynamic constructor so that the static tree may thereafter be updated dynamically via insertion or deletion of individual tuples.
The Java and C++ implementations are also available at the following URL.